An electron accelerated by 200 V, enters in magnetic field. If its velocity is 100 m/s, then (e/m) for it will be.
(A) 1.75 * 10^10
(B) 1.75 * 10^11
(C) 1.75 * 10^9
(D) 1.75 * 10^6

Grade:12th pass

2 Answers

Vikas TU
5 years ago
The kinetic energy would be after accelerateed by 200 V would be:
0.5*mv^2 = eV
0.5*v^2/V = e/m
e/m = (0.5*100^2)/200 => 25.
 
Asha
2 years ago
According to the measurement by Thomson, the value of e/m is 1.77*10^11 C/kg. So without calculation we can say the value